# Adapting Large Multimodal Models to Distribution Shifts: The Role of In-Context Learning

### v1

@@ Abstract
Recent studies indicate that large multimodal models (LMMs) are highly robust against natural distribution shifts, often surpassing previous baselines. Despite this, domain-specific adaptation is still necessary, particularly in specialized areas like healthcare. Due to the impracticality of fine-tuning LMMs given their vast parameter space, this work investigates in-context learning (ICL) as an effective alternative for enhancing LMMs’ adaptability. We find that the success of ICL heavily relies on the choice of demonstration, mirroring challenges seen in large language models but introducing unique complexities for LMMs facing distribution shifts. Our study addresses this by evaluating an unsupervised ICL method, TopKNearestPR, which selects in-context examples through a nearest example search based on feature similarity. We uncover that its effectiveness is limited by the deficiencies of pre-trained vision encoders under distribution shift scenarios, evidenced by their zero-shot capabilities barely outperforming random guesses. To address these challenges, we propose InvariantSelectPR, a novel method leveraging Class-conditioned Contrastive Invariance (CCI) for more robust demonstration selection. Specifically, CCI enhances pre-trained vision encoders by improving their discriminative capabilities across different classes and ensuring invariance to domain-specific variations. This enhancement allows the encoders to effectively identify and retrieve the most informative examples, which are then used to guide LMMs in adapting to new query samples under varying distributions. Our experiments show that InvariantSelectPR substantially improves the adaptability of LMMs, achieving significant performance gains on benchmark datasets, with a $34.2\%$$\uparrow$ accuracy increase in 7-shot on Camelyon17 and $16.9\%$$\uparrow$ increase in 7-shot on HAM10000 compared to the baseline zero-shot performance. Our code will be publicly available at: https://github.com/jameszhou-gl/icl-distribution-shift.
$\dagger$\$$\dagger$\$footnotetext: These authors contributed equally to this work.
@@
Figure 1: Comparative illustration of (a) zero-shot transfer, which relies on LMMs’ pre-trained knowledge to respond to queries, potentially leading to a large distribution gap, and (b) in-context learning (ICL), which introduces an example from a closer distribution to bridge this gap. This work investigates different retrieval methods for selecting effective ICL examples. (c) The efficacy of one-shot ICL is showcased in guiding LMMs with shifted distributions.
1 Introduction
@@
Machine learning models are essential in areas such as climate modeling, biomedicine, and autonomous driving, where they need to reliably manage deviations from their training data known as distribution shifts (Park et al., 2021; Zhou et al., 2022, 2023b). Traditional methods like domain adaptation and domain generalization have been somewhat effective but still fall short in addressing these shifts, as confirmed by several empirical studies (Gulrajani and Lopez-Paz, 2020; Wiles et al., 2022). However, the emergence of foundation models, characterized by their extensive and diverse pretraining, offers new possibilities for enhancing adaptability to these challenges (Bommasani et al., 2021; Radford et al., 2021; Shu et al., 2023). Specifically, large multimodal models (LMMs) (Yang et al., 2023b) such as GPT-4V (OpenAI, 2023), and Gemini (Team et al., 2023) have shown superior adaptability. Their zero-shot [^1] We adopt the zero-shot setting in the CLIP study (Radford et al., 2021), which enables zero-shot transfer via paired image-text features without extra linear probes. It differs from the classical zero-shot of generalizing to unseen categories (Chen et al., 2021, 2022, 2023). capabilities have been found to frequently outperform the performance of traditional fine-tuned models in natural datasets (Han et al., 2024).
@@
Despite recent advances, domain-specific adaptation remains a significant challenge, especially in healthcare and scientific research (Han et al., 2024). While LMMs like Google DeepMind’s Med-Gemini offer fine-tuned versions for medical tasks (Saab et al., 2024), their block-box nature and massive parameter sets make traditional fine-tuning impractical for researchers without extensive computational resources. This highlights the urgent need for more feasible adaptation techniques. In-context learning (ICL), which allows models to adapt during inference without parameter adjustments, emerges as a promising alternative (Brown et al., 2020; Liu et al., 2021; Min et al., 2022; Dong et al., 2022). While the effectiveness of ICL is recognized within large language models (LLMs), its application for improving adaptability in LMMs under distribution shifts is less explored.
@@
As depicted in Figure 1, we hypothesize that equipping LMMs with context examples that include task-specific information and details about the query sample can substantially enhance their performance. Our research starts with a thorough evaluation of ICL’s capacity to tailor LMMs (§2) to specific domains, particularly healthcare research, where there is a clear necessity for domain-specific fine-tuning (Han et al., 2024). We discover that the success of ICL heavily depends on the choice of demonstrations, supporting findings from previous LLMs studies (Liu et al., 2021; Min et al., 2022; Dong et al., 2022). Although the significant impact of demonstration selection on ICL performance is not entirely unexpected, the challenge of selecting demonstrations for LMMs under distribution shifts remains unexplored. To address this, we re-examine the unsupervised retrieval of in-context examples (§3.1), TopKNearestPR, traditionally used in LLMs. This method, discussed in (Liu et al., 2021; Zhang et al., 2024), uses feature similarity to pinpoint contextually relevant ICL examples.
@@
However, the TopKNearestPR approach faces considerable challenges when applied to LMMs during distribution shifts. Notably, using pretrained vision encoders like CLIP-ViT [^2] https://huggingface.co/openai/clip-vit-base-patch16, zero-shot performance often remains at levels comparable to random guessing in specialized domains, as shown in a recent study (Han et al., 2024). This poor performance reveals a critical limitation in these encoders: they struggle to recognize and adapt to the subtle variations in new distributions, which compromises the reliability of visual feature similarities for selecting effective demonstrations. To tackle these challenges, we propose InvariantSelectPR, a novel method designed specifically for scenarios involving distribution shifts (§3.2). This approach employs Class-conditioned Contrastive Invariance (CCI) to choose demonstrations based on domain-invariant features, which are inherently robust to distributional changes (Zhou et al., 2024). This retriever method is distinctively crafted for distribution shifts, ensuring the resilience of selected in-context examples in varying conditions. Our empirical results demonstrate that InvariantSelectPR significantly improves the adaptability of LMMs, achieving notable accuracy improvements, i.e., a 34.2% accuracy improvement in 7-shot on Camelyon17 and a 16.9% accuracy increase in 7-shot on HAM10000 over the zero-shot baseline.
@@
Our contributions and the key findings are summarized as follows: (1) To the best of our knowledge, this work takes the first step towards deeply understanding the role of in-context learning as an effective strategy for enhancing the adaptability of LMMs under distribution shifts (§2). (2) We introduce InvariantSelectPR, a novel in-context retrieval framework specifically developed to tackle distribution shifts (§3). (3) Through extensive experiments on four benchmark datasets (§4), our InvariantSelectPR method shows substantial enhancements over zero-shot generalization capabilities.
@@
Figure 2: Motivation illustration: (a) Performance comparison between Zero-shot and RandomPR, illustrating the limitations of random in-context example selection across four datasets, where one-shot RandomPR often underperforms compared to zero-shot. (b) Analysis of 77 query samples from the target domain, hospital_3 in Camelyon17, using 50 distinct one-shot examples to examine performance variability. Mean values are marked in blue, and variance is represented by black lines, highlighting the significant impact of example selection on model accuracy. If appropriate in-context samples are chosen, there is a potential for gains up to 40.25%.
2 Motivation: ICL Demonstrations under Distribution Shifts
@@
In this section, we evaluate ICL’s capability to enhance LMM adaptability. Starting with RandomPR, we randomly select in-context examples from source domain data without relevance to the target task. Our evaluation uses the Gemini model, noted for its zero-shot capabilities, across four medical datasets typically needing domain-specific fine-tuning (Han et al., 2024). We compare one-shot RandomPR with the zero-shot baseline for a preliminary investigation. According to Figure 2(a), while RandomPR presents a slight decrease of 0.3% on the Camelyon17 dataset, it leads to a substantial performance decline of 4.2%, 3.7%, and even 8.2% on the HAM10000, NIH_Chest, and COVID datasets respectively. Despite its conceptual simplicity, our empirical results suggest that random in-context example selection often fails to fulfill the essential requirement for effective model adaptation—providing informative and contextually appropriate demonstrations.
@@ 3 Methodology
3.1 TopKNearestPR: Enhancing Context Relevance
@@
3.2 InvariantSelectPR: Tailored for Distribution Shift Adaptation
@@
Facilitating Class-conditioned Contrastive Invariance.  InvariantSelectPR is centered around the Class-conditioned Contrastive Invariance (CCI) mechanism, which aims to improve the model’s ability to distinguish between classes while maintaining stability across domain-specific variations (Zhou et al., 2024). This is achieved by promoting similarity among instances of the same class from different domains and highlighting differences between classes. Using the class token embedding [CLS], $\bm{x}_{N}$, from the final vision transformer (ViT) layer, the CCI loss is defined as:
@@
Here, $\bm{z}_{N^{\prime}}$ is a positive sample of $\bm{z}_{N}$ from the same class but possibly a different domain, and $\bm{z}_{k}$ signifies a negative sample of $\bm{z}_{N}$ from a different class. $\tau$ denotes the temperature parameter in contrastive learning (Chen et al., 2020; Zhou et al., 2023a). This formulation ensures that the learned representations are both discriminative and invariant, crucial for adapting to new distributions.
@@
In-Context Selection Through Enhanced Invariance.  After fine-tuning the vision encoder with the combined losses in Eq. (5), we leverage refined features to assess the similarity between the target samples and in-context examples. By ensuring these similarities reflect both visual resemblance and domain invariance, the k-shot examples with the highest similarity scores are then selected.
4 Experiments
Datasets Overview. We use four benchmark datasets to explore distribution shifts, particularly emphasizing domain-specific fine-tuning (Han et al., 2024). Camelyon17 (Bandi et al., 2018) features 450,000 patches from breast cancer images across five hospitals. HAM10000 (Tschandl et al., 2018) offers dermatoscopic images critical for skin cancer detection. The NIH_Chest dataset (Wang et al., 2017) includes over 112,000 X-ray images annotated for thoracic diseases. The COVID dataset (Han et al., 2021) provides diverse pneumonia detection data, including COVID-19 cases, from various hospitals. We analyze a practical subset, random_1, with 450 samples [^3] Available at https://github.com/jameszhou-gl/gpt-4v-distribution-shift.
Implementation Details. We compare three retrieval methods—RandomPR, TopKNearestPR and InvariantSelectPR—against the baseline zero-shot capability. We employ vit_large_patch14_224_clip_laion2b configuration from the timm library, exploring variations in backbone configurations further in §4.3.2. The Gemini model is employed as the primary LMM due to its superior zero-shot performance across varied datasets (Han et al., 2024) and its stable log-linear improvement in performance with an increasing number of ICL examples, as observed in a concurrent study (Jiang et al., 2024). Our main results (§4.1) focus on one-shot performance, with additional insights on the impact of different numbers of shots in §4.3.3. We also include other leading LMMs, such as GPT-4V and Claude (Anthropic, 2023), in our extended analysis in §4.3.4.
[…]
4.1 Main Results
In Tables 1 and 2, our analysis of four benchmark datasets provides a detailed examination of how different ICL methods perform under distribution shifts. The zero-shot approach highlights the inherent ability of LMMs to adapt to new domains without retraining. However, the effectiveness of this adaptability varies significantly with different ICL strategies. The RandomPR strategy, which employs a stochastic method for selecting in-context examples, yields inconsistent results. For instance, on the Camelyon17 dataset, it leads to a slight decrease of 0.3% in accuracy, but it largely underperforms on the COVID, HAM10000, and NIH_Chest datasets, with accuracy decreases of 8.2%, 4.2%, and 3.7%, respectively. This highlights the unpredictable performance of RandomPR across different conditions. Conversely, TopKNearestPR uses a pre-trained vision encoder to identify feature similarities for example selection, leading to a 5.74% improvement on the Camelyon17, which demonstrates the benefits of a more targeted approach in example selection. Despite this success, the method sees declines of 3.9% and 2.9% on the HAM10000 and NIH_Chest datasets, respectively, indicating a lack of consistent performance across all test scenarios. The most effective strategy, InvariantSelectPR, consistently outperforms other methods, significantly exceeding the zero-shot baseline across all datasets, especially achieving remarkable gains of 8.3% on HAM10000 and 8.6% on Camelyon17. These results underscore the importance of advanced in-context example selection techniques in adapting LMMs to distribution shifts. Despite notable gains, the improvements with InvariantSelectPR on NIH_Chest and COVID are modest. In Figure 4, the incremental improvements by InvariantSelectPR align with those from fine-tuned encoders, which generally surpass the fine-tuning approach by 1% to 6%. This suggests that when fine-tuning itself is minimally effective, ICL strategies yield limited enhancements. Future research thus focuses on more sophisticated methods to enhance invariance, beyond fundamental domain-invariance in this work.
Table 1: Performance comparison of three retrieval methods against the zero-shot approach, illustrating accuracy improvements or decreases on Camelyon17 and COVID datasets. Mean and standard deviation values are calculated over three independent runs, with the best results highlighted in bold.

### v2

@@ Abstract
Recent studies indicate that large multimodal models (LMMs) potentially act as general-purpose assistants and are highly robust against different distributions. Despite this, domain-specific adaptation is still necessary particularly in specialized areas like healthcare. Due to the impracticality of fine-tuning LMMs given their vast parameter space, this work investigates in-context learning (ICL) as an effective alternative for enhancing LMMs’ adaptability. Our study proceeds this by evaluating an unsupervised ICL method which selects in-context examples through a nearest example search based on feature similarity. We uncover that its effectiveness is limited by the deficiencies of pre-trained vision encoders under distribution shift scenarios, evidenced by their zero-shot capabilities barely outperforming random guesses. To address these challenges, we propose InvariantSelectPR, a novel method leveraging Class-conditioned Contrastive Invariance (CCI) for more robust demonstration selection. Specifically, CCI enhances pre-trained vision encoders by improving their discriminative capabilities across different classes and ensuring invariance to domain-specific variations. This enhancement allows the encoders to effectively identify and retrieve the most informative examples, which are then used to guide LMMs in adapting to new query samples under varying distributions. Our experiments show that InvariantSelectPR substantially improves the adaptability of LMMs, achieving significant performance gains on benchmark datasets, with a $34.2\%$$\uparrow$ accuracy increase in 7-shot on Camelyon17 and $16.9\%$$\uparrow$ increase in 7-shot on HAM10000 compared to the baseline zero-shot performance. Our code will be publicly available at: https://github.com/jameszhou-gl/icl-distribution-shift.
Index Terms:
Large multimodal models, Distribution shifts, In-context learning.
@@
Figure 1: Comparative illustration of (a) zero-shot transfer, which relies on LMMs’ pre-trained knowledge to respond to queries, potentially leading to a large distribution gap between pre-training data and query samples, and (b) in-context learning (ICL), which introduces an example from a closer distribution with query sample to bridge this gap. This work investigates different retrieval methods for selecting effective ICL examples.
I Introduction
@@
Machine learning models are essential in areas such as climate modeling, biomedicine, and autonomous driving, where they need to reliably manage deviations from their training data known as distribution shifts [1, 2, 3]. Traditional methods like domain adaptation (DA) and domain generalization (DG) have been somewhat effective but still fall short in addressing these shifts, as confirmed by several empirical studies [4, 5]. However, the emergence of foundation models, characterized by their extensive and diverse pretraining, offers new possibilities for enhancing adaptability to these challenges [6, 7, 8]. Specifically, large multimodal models (LMMs) [9] such as GPT-4V [10], Claude [11] and Gemini [12] have shown superior adaptability. Their zero-shot [^1] In this study, zero-shot refers to the ability to apply models to new tasks without additional training [7]. It differs from traditional zero-shot learning of generalizing to unseen categories [13, 14]. capabilities have been found to frequently outperform the performance of traditional fine-tuned models in natural datasets [15].
@@
Despite recent advances, domain-specific adaptation remains a significant challenge, especially in healthcare [15]. While LMMs like Google DeepMind’s Med-Gemini offer fine-tuned versions for medical tasks [16], their block-box nature and massive parameter sets make traditional fine-tuning impractical for researchers without extensive computational resources. This highlights the urgent need for more feasible adaptation techniques. In-context learning (ICL), which allows models to adapt during inference without parameter adjustments, emerges as a promising alternative [17, 18, 19, 20]. While the effectiveness of ICL is recognized within large language models (LLMs), its application for improving adaptability in LMMs under distribution shifts is less explored.
@@
As depicted in Figure 1, we hypothesize that equipping LMMs with context examples that include task-specific information and details about the query sample can substantially enhance their performance. Our research starts with a thorough evaluation of ICL’s capacity to tailor LMMs (§III-B) to specific domains, particularly healthcare research, where there is potentially large distribution shift through proxy measures (§III-A). We discover that the success of ICL heavily depends on the choice of demonstrations. To address this, we re-examine the unsupervised retrieval of in-context examples (§IV-A), TopKNearestPR, traditionally used in LLMs. This intuitive method uses feature similarity to pinpoint contextually relevant ICL examples [18, 21].
@@
However, in scenarios of distribution shifts, the TopKNearestPR approach faces considerable challenges when applied to LMMs. Notably, using pretrained vision encoders like CLIP-ViT [^2] https://huggingface.co/openai/clip-vit-base-patch16, zero-shot performance often remains at levels comparable to random guessing in specialized domains, as shown in Figure 2. This poor performance reveals a critical limitation in these encoders: they struggle to recognize and adapt to the subtle variations in new distributions, which compromises the reliability of visual feature similarities for selecting effective demonstrations. To tackle these challenges, we propose InvariantSelectPR, a novel method designed specifically for scenarios involving distribution shifts (§IV-B). This approach employs Class-conditioned Contrastive Invariance (CCI) to choose demonstrations based on domain-invariant features, which are inherently robust to distributional changes [22]. This retriever method is distinctively crafted for distribution shifts, ensuring the resilience of selected in-context examples in varying conditions. Our empirical results demonstrate that InvariantSelectPR significantly improves the adaptability of LMMs, achieving notable accuracy improvements, i.e., a 34.2% accuracy improvement in 7-shot on Camelyon17 and a 16.9% accuracy increase in 7-shot on HAM10000 over the zero-shot baseline.
@@
Our contributions and the key findings are summarized as follows: (1) To the best of our knowledge, this work takes a first step towards deeply understanding in-context learning as an effective strategy for adapting LMMs to distribution shifts (§III). (2) We introduce InvariantSelectPR, a novel in-context retrieval framework specifically developed to tackle distribution shifts (§IV). (3) Through extensive experiments on four benchmark datasets (§V), our InvariantSelectPR method shows substantial enhancements over LMMs’ zero-shot capabilities.
Figure 2: A proxy task to evaluate potential distribution shifts in LMMs, illustrating zero-shot performance across various datasets compared to random guessing. Red horizontal lines indicate the average performance of LMMs for each dataset, and their minor deviations from random guessing highlight significant shifts, particularly in medical contexts such as Camelyon17, HAM10000, NIH-Chest, and COVID datasets.
II Related Work
II-A Distribution Shifts
The literature on distribution shifts categorizes mitigation approaches into two primary strategies: domain adaptation and domain generalization. Domain adaptation techniques, well-established for scenarios where the target domain is known during training, recalibrate models according to the target data’s statistical properties [23]. These techniques encompass deep transfer learning, which aligns feature distributions between source and target domains [24], unsupervised methods that minimize domain discrepancies [25]. In contrast, domain generalization addresses the more daunting challenge of excelling in completely unseen domains. Strategies here include aligning features across multiple source domains [26], separating domain-specific from domain-general features [27, 28], employing meta-learning for optimization across various domains [29, 30], and using data augmentation to mimic domain variability [31, 32]. Recent studies have observed that LMMs have demonstrated exceptional adaptability when dealing with natural datasets but cannot handle the distributions in specialized areas such as healthcare [7, 33]. This observation motivates this paper’s to explore distribution shifts in LMMs through proxy measures and investigate the ICL strategies under distribution shifts.
II-B In-Context Learning
In-context learning (ICL), particularly defined in GPT-3 [17], originated in LLMs for natural language processing (NLP) tasks. ICL is a proven effective paradigm that leverages context augmented with a few examples to enable LLMs to make predictions [20, 34, 35, 19, 20, 36, 37, 38, 39]. The choice of these in-context examples critically impacts performance, as evidenced by studies demonstrating that selecting nearest neighbors based on sentence encoders can significantly enhance the few-shot capabilities of models like GPT-3 [18, 40, 41]. While ICL is established in NLP, it is emerging in visual and multimodal LLMs. The study Flamingo [42] marks the earliest exploration of visual ICL, with subsequent studies validating the importance of example selection in image painting models [43, 44, 21], visual understanding [45, 46], and diffusion models [47]. Unlike prior work, this paper uniquely focuses on deeply understanding the role of ICL under distribution shifts, taking a first step in this direction.
@@ II-B In-Context Learning
Figure 3: ICL Demonstrations under Distribution Shifts: (a) Performance comparison between Zero-shot and RandomPR, illustrating the limitations of random in-context example selection across four datasets, where one-shot RandomPR often underperforms compared to zero-shot. (b) Analysis of 77 query samples from the target domain, hospital_3 in Camelyon17, using 50 distinct one-shot examples to examine performance variability. Mean values are marked in blue, and variance is represented by black lines, highlighting the significant impact of example selection on model accuracy. If appropriate in-context samples are chosen, there is a potential for gains up to 40.25%.
III Motivation
III-A Evaluating Distribution Shifts in LMMs Through Proxy Measures
Distribution shifts traditionally refer to discrepancies between training and test data distributions [48, 49]. Traditional domain adaptation (DA) and domain generalization (DG) tackle these shifts by fine-tuning models on training data to enhance performance across various yet related test distributions. Instead, LMMs are engineered to serve as general-purpose assistants [50], effectively operating in an efficient way in zero-shot or few-shot modes without the need for parameter adjustments. Therefore, we conceptualize distribution shifts in LMMs as the gap between the data distributions seen during their pre-training and the data distributions in test scenarios. This shift is particularly challenging to quantify directly because details of the pre-training data are often not publicly available.
To tackle this, we adopt using a proxy task approach, inspired by similar strategies from related research on data contamination [51, 52]. This involves comparing the zero-shot performance against a random guessing baseline to infer potential distribution shifts. Minor deviations from this baseline indicate a potential distribution shift. When test data features are not well-represented within the LMM’s pre-training data, we expect the model to perform only slightly better than random guessing. Such small performance differences indicate that the model is facing unfamiliar data, highlight a substantial shift in distribution.
To build a suitable benchmark for our study, we initially collected data from eleven public datasets. We then applied our proxy method to assess distribution differences, specifically, we evaluated four LMMs across various datasets and measured their zero-shot effectiveness, and calculated these results against the expected performance of random guessing [^3] We follow the procedures in https://github.com/jameszhou-gl/gpt-4v-distribution-shift, as illustrated in Figure 2. Our findings reveal pronounced shifts in medical datasets compared to natural datasets, such as Camelyon17, HAM10000, NIH-Chest, and COVID. These four datasets serve as our primary benchmarks. This analysis not only underscores the distribution shifts but also highlights the necessity of adaptive strategies for LMMs, which is the reason that we investigate and design effective ICL methods.
III-B ICL Demonstrations under Distribution Shifts
@@
In this section, we evaluate ICL’s capability to enhance LMM adaptability. Starting with RandomPR, we randomly select in-context examples from source domain data without relevance to the target task. Our evaluation uses the Gemini model, noted for its zero-shot capabilities, across four medical datasets typically needing domain-specific fine-tuning [15]. We compare one-shot RandomPR with the zero-shot baseline for a preliminary investigation. According to Figure 3(a), while RandomPR presents a slight decrease of 0.3% on the Camelyon17 dataset, it leads to a substantial performance decline of 4.2%, 3.7%, and even 8.2% on the HAM10000, NIH_Chest, and COVID datasets respectively. Despite its conceptual simplicity, our empirical results suggest that random in-context example selection often fails to fulfill the essential requirement for effective model adaptation—providing informative and contextually appropriate demonstrations.
@@ IV Methodology
IV-A TopKNearestPR: Enhancing Context Relevance
@@
IV-B InvariantSelectPR: Tailored for Distribution Shift Adaptation
@@
IV-B1 Facilitating Class-conditioned Contrastive Invariance
InvariantSelectPR is centered around the Class-conditioned Contrastive Invariance (CCI) mechanism, which aims to improve the model’s ability to distinguish between classes while maintaining stability across domain-specific variations [22]. This is achieved by promoting similarity among instances of the same class from different domains and highlighting differences between classes. Using the class token embedding [CLS], $\bm{x}_{N}$, from the final vision transformer (ViT) layer, the CCI loss is defined as:
@@ IV-B1 Facilitating Class-conditioned Contrastive Invariance
Here, $\bm{z}_{N^{\prime}}$ is a positive sample of $\bm{z}_{N}$ from the same class but possibly a different domain, and $\bm{z}_{k}$ signifies a negative sample of $\bm{z}_{N}$ from a different class. $\tau$ denotes the temperature parameter in contrastive learning [53, 54]. This formulation ensures that the learned representations are both discriminative and invariant, crucial for adapting to new distributions.
@@
IV-B2 In-Context Selection Through Enhanced Invariance
After fine-tuning the vision encoder with the combined losses in Eq. (5), we leverage refined features to assess the similarity between the target samples and in-context examples. By ensuring these similarities reflect both visual resemblance and domain invariance, the k-shot examples with the highest similarity scores are then selected.
TABLE I: Detailed statistics of the datasets used in the experiments.
Dataset Prediction Task # Domains # Classes Example Image
Camelyon17 Tumor detection 5 2 [Uncaptioned image]
HAM10000 Skin disease classification 4 7 [Uncaptioned image]
NIH_Chest Lung disease diagnosis 2 15 [Uncaptioned image]
COVID Pneumonia type classification 2 3 [Uncaptioned image]
TABLE II: Performance comparison of three retrieval methods against the zero-shot approach, illustrating accuracy improvements or decreases on Camelyon17 and COVID datasets. The best and second-best results are marked in Red and Blue respectively. Mean and standard deviation values are calculated over three independent runs. $\rightarrow$” denotes the test scenario.